### v1

# EXCGEC: A Benchmark of Edit-wise Explainable Chinese Grammatical Error Correction

Abstract
Existing studies explore the explainability of Grammatical Error Correction (GEC) in a limited scenario, where they ignore the interaction between corrections and explanations. To bridge the gap, this paper introduces the task of EXplainable GEC (EXGEC), which focuses on the integral role of both correction and explanation tasks. To facilitate the task, we propose EXCGEC, a tailored benchmark for Chinese EXGEC consisting of 8,216 explanation-augmented samples featuring the design of hybrid edit-wise explanations. We benchmark several series of LLMs in multiple settings, covering post-explaining and pre-explaining. To promote the development of the task, we introduce a comprehensive suite of automatic metrics and conduct human evaluation experiments to demonstrate the human consistency of the automatic metrics for free-text explanations. [^1] All the codes and data will be released after the review.
{CJK*}
UTF8gbsn
EXCGEC: A Benchmark of Edit-wise Explainable Chinese
Grammatical Error Correction
Jingheng Ye1thanks: indicates equal contribution., Shang Qin1∗, Yinghui Li1, Xuxin Cheng2, Libo Qin3 Hai-Tao Zheng1,4thanks: Corresponding author: Hai-Tao Zheng. (E-mail: zheng.haitao@sz.tsinghua.edu.cn), Peng Xing1, Zishan Xu1, Guo Cheng1, Zhao Wei5 1Tsinghua University, 2Peking University 3Central South University, 4Peng Cheng Laboratory, 5Tencent {yejh22,qin-s23}@mails.tsinghua.edu.cn
1 Introduction
Despite the notable advancements in Grammatical Error Correction (GEC) Bryant et al. (2023); Zhao et al. (2018); Ma et al. (2022); Ye et al. (2023a), there still exists a lack of profound examination into the explainability of GEC Dwivedi et al. (2023), which is critical in educational scenarios for L2 (Language second)-speakers Wang et al. (2021) or school-age children Li et al. (2023c). These mainstream users, who often face challenges in creating grammatically accurate and fluent texts, may be confused or even misguided if they are provided with limited access to only corrective texts. Therefore, augmenting the explainability of GEC is unquestionably beneficial for the progression of the GEC community as well as related fields, such as essay scoring Li et al. (2022b); Stahl et al. (2024), intelligent tutoring systems Montenegro-Rueda et al. (2023) and other emerging educational scenarios Li et al. (2024c); Du et al. (2024).
Figure 1: Task definitions of GEC, GEE and EXGEC. For the error description of EXGEC, we highlight evidence words, linguistic knowledge, error causes, and revision advice parts in different colors.
As illustrated in Figure 1, existing tasks like GEC and Grammatical Error Explanation (GEE) typically address either correction or explanation, ignoring the interaction between the two. To bridge the gap, we introduce the task of EXplainable Grammatical Error Correction (EXGEC). By integrating these two tasks, EXGEC enables systems to elucidate the linguistic knowledge and reasoning mechanism underlying predicted corrections, thereby achieving the best of both worlds. Additionally, EXGEC can function as a test bed for determining the explainable abilities of large language models (LLMs) and identifying any unintended biases and risks in educational scenarios.
To facilitate EXGEC, we present EXCGEC, a tailored benchmark for Chinese EXGEC, featuring the design of hybrid edit-wise explanations. Each explanation, based on a particular edit, consists of three elements: 1) Error types, which allow learners to absorb syntax and semantic knowledge in an inductive way Fei et al. (2023). We establish a hierarchical and pragmatic two-tier taxonomy for Chinese grammatical errors. 2) Error severity levels ranging from 1 $\sim$ 5 points, which are beneficial to prioritize core corrections. 3) Error descriptions, presented as the form of natural language explanation Camburu et al. (2018); He et al. (2023), provide evidence words, relevant linguistic knowledge or syntax rules, error causes, and revision advice for edits. The design provides more detailed and faithful guidance for learners, allowing them to comprehend each grammatical error committed. This is unlikely achievable for other designs such as example-based Kaneko et al. (2022) or sentence-level explanations Nagata et al. (2021).
Stimulated by recent success of synthetic data generation Shum et al. (2023); Whitehouse et al. (2023), we employ a semi-automatic dataset construction solution to enhance efficiency, while minimising annotation costs. Initially, we synthesize the EXCGEC dataset by prompting GPT-4 Liu et al. (2024). Then we hire native annotators to filter invalid data and provide a detailed analysis of invalid data, ensuring the quality of the dataset Ding et al. (2024). We finally obtain 8,216 clean explanation-augmented samples for benchmarking. We also introduce automatic metrics to evaluate performance across both tasks, and conduct human evaluation experiments to ascertain the correlation between these metrics and human judgement, thus demonstrating their effectiveness.
Based on the benchmark, we develop EXGEC baseline models that can perform both the correction and explanation tasks in either post-explaining (correct-then-explain) or pre-explaining (explain-then-correct) sequences. Particularly, we design Correct-Then-Explain (COTE) decoding algorithm for post-explaining models. Benchmarking various series of open-source LLMs has yielded several intriguing findings. For example, post-explaining models display higher performance than pre-explaining models. Moreover, COTE significantly enhances performance by alleviating the alignment workload for the LLMs. Our contributions in this paper are listed as follows:
  • (1)
    We introduce the EXGEC task and establish a corresponding benchmark consisting of a Chinese EXGEC dataset and a comprehensive set of automatic metrics, contributing to the stable development of the field of EXGEC.
  • (2)
    We develop EXGEC baseline models and investigate the abilities of various LLMs using our proposed benchmark.
  • (3)
    We conduct detailed analyses on our proposed dataset and baselines to gain further insights. Human evaluation experiments are also conducted to confirm the effectiveness of automatic metrics for error descriptions.
2 Related Work
Exploration of explainable GEC has witnessed a paradigm shifting from fine-tuning Yang et al. (2020); Ye et al. (2022); Li et al. (2022c); Ye et al. (2023b); Kaneko and Okazaki (2023); Huang et al. (2023) to prompting Li et al. (2024b); Tan et al. (2023); Yu et al. (2024); Zhao et al. (2024); Xu et al. (2024); Qin et al. (2024a, b), with the focus being local explanations of individual predictions. Fei et al. (2023) construct an explainable GEC dataset EXPECT, which is annotated with evidence words and error types based on the standard GEC benchmark Bryant et al. (2019). However, EXPECT falls short of flexibility due to the lack of natural language explanations. To fill the gap, Song et al. (2023) propose the task of grammatical error explanation. They observe that GPT-4 suffers from identifying and explaining errors with limited access to only parallel source-target pairs. To address this issue, they fine-tune an extra LLM as an edit extractor, which is trained on synthesized data. However, all these studies overlook the benefits of effectiveness and efficiency brought by multi-task learning both correction and explanation tasks, which is extensively explored in this work.
On the other hand, a similar task, called the feedback comment generation task, focuses on sentence-level explanations Liu et al. (2022); Dong et al. (2023); Sun et al. (2022); Li et al. (2023b); Huang et al. (2024). However, it suffers from high cost associated with data annotation Nagata et al. (2020). Furthermore, due to the complexity of the task Nagata (2019), it is often explored with limited access to only a subset of grammatical error types.
3 Task Definition
3.1 Grammatical Error Correction
GEC has been studied for decades, witnessing the shift from rule-based methods to LLM-based methods. Formally, given an ungrammatical text (source text) $X=\{x_{1},x_{2},\cdots,x_{T}\}$, a GEC model is required to correct $X$ into a grammatically correct counterpart (target text) $Y=\{y_{1},y_{2},\cdots,y_{T^{\prime}}\}$ without changing the original semantic as far as possible. Typically, GEC is usually treated as a sequence-to-sequence (Seq2Seq) task, the training objective of which is formulated as follows:
$\mathcal{L}_{\operatorname{GEC}}=-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X)$ (1)
3.2 Grammatical Error Explanation
GEE has been explored in several methodologies, including sentence-level explanation and edit-wise explanation. Since sentence-level explanations suffer from over-generalization and confusion especially when a sentence contains multiple grammatical errors, this work focuses solely on edit-wise explanations. Given a source text $X$ and its target counterpart $Y$, the GEE model needs to explain each grammatical error $e_{i}$ in $X$. Specifically, GEE is typically solved in a two-step pipeline consisting of edit extraction and edit-wise explanation. 1) Edit extraction produces an edit set $E=\{e_{1},e_{2},\cdots,e_{n}\}$ that represent grammatical errors in $X$ and also clarify the transformation from ungrammatical segments of $X$ to target segments of $Y$. Typically, an edit contains four key elements: source position $sp$, source content $sc$, target position $tp$, and target content $tc$. The process of edit extraction can be easily accomplished using alignment-based evaluation toolkits like ERRANT Bryant et al. (2017); Felice et al. (2016) and CLEME Ye et al. (2023c). 2) Edit-wise explanation generates a set of explanations $E^{\prime}=\{e^{\prime}_{1},e^{\prime}_{2},\cdots,e^{\prime}_{n}\}$, with each explanation $e^{\prime}_{i}$ corresponding to $e_{i}$, given the source and the target texts. Although the design of explanation varies across related work Song et al. (2023); Zhao et al. (2024), the typical training objective of GEE models is presented as follows:
$E=f(X,Y)$ (2)
$\mathcal{L}_{\operatorname{GEE}}=-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i})$ (3)
where $f:(X,Y)\to E=\{(sp_{i},sc_{i},tp_{i},tc_{i})\}^{n}_{i=1}$ is the edit extraction function used to extract edits of $X$ and $Y$, and $n$ is the number of edits.
Existing studies Song et al. (2023); Fei et al. (2023) focus on developing GEE models that can generate more reasonable explanations. However, an extra GEC model is compulsory to allow GEE models to generate explanations if only source texts are offered, thus resulting in an issue of low efficiency. Furthermore, there exists a gap between GEC and GEE models if they are trained on different data with domain shifts.
Figure 2: Overview of benchmark construction and model development. We show the inference process of the post-explaining model in particular.
3.3 Explainable Grammatical Error Correction
To get rid of the drawbacks brought by the nature of GEE, we propose the EXGEC task which aims to perform both correction and explanation tasks simultaneously. The motivation for combining these two tasks majorly falls on two aspects. First, a branch of existing studies Wiegreffe and Marasovic (2021); Hartmann and Sonntag (2022); Li et al. (2022a, 2024a) have demonstrated training with access to human explanations can improve model performance. It is also intuitive that either of the GEC and GEE tasks can mutually benefit from each other when training in a multi-task manner. Second, it is more time-saving and cost-efficient to deploy a single EXGEC model rather than two detached models in foreign language education platforms.
In this task, the only input element is an ungrammatical source text $X$, and the EXGEC model learns to output both the grammatical target text $Y$ and explanations $E^{\prime}$. Similar to GEE, EXGEC follows the edit-wise style of explanation, and it is categorized into two different settings by the order of correction and explanation tasks, with the basic scheme of multi-task learning.
Post-explaining.
Models are trained first to generate target texts Camburu et al. (2018), which allows the explanations to be explicitly conditioned on the target texts, thus ensuring high faithfulness of explanations towards the target texts. The training objective is as follows:
$\mathcal{L}_{\operatorname{post}}=$ $-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X)$ (4)
$-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i})$
The inference of post-explaining models is represented as follows:
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{post}}(X)$ (5)
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{post}}(X,Y,f(X,\hat{Y}))$ (6)
With the target texts generated ahead, post-explaining models can output explanations conditioned on the specific edits that are extracted by aligning the source and the target texts, thus improving the accuracy and faithfulness of explanations.
Pre-explaining.
This type of models are trained conversely, whose mechanism is similar to the Chain of Thought (CoT) technique Chai et al. (2024). Pre-explaining models are supposed to make full use of synthesized explanations to generate elaborated target texts. With minimal modification from Equation (4), the training objective of pre-explaining models is as follows:
$\mathcal{L}_{\operatorname{pre}}=$ $-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X)$ (7)
$-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X,E^{\prime})$
Notably, pre-explaining models may struggle to generate well-formed edit-wise explanations due to the inaccessibility to the edit extraction function $f$, which necessitates both the source and the target texts. Similarly, the inference of pre-explaining models is presented as follows:
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{pre}}(X)$ (8)
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{pre}}(X,E^{\prime})$ (9)
4 EXCGEC Benchmark
To facilitate the development of EXGEC task, we construct EXCGEC, the first benchmark for explainable Chinese GEC particularly. As illustrated in Figure 2, we begin by the process of data curation, which consists of explanation design in Section 4.1, explanation synthesis and refinement in Section 4.2. Then we gain an in-depth understanding of GPT-4 Achiam et al. (2023) in EXGEC by further analyzing generated data in Section 4.3, where we summarize common failure modes in invalid instances. Finally, we introduce a series of automatic metrics for evaluating explanations in Section 4.4.
4.1 Explanation Design
In the pursuit of comprehensiveness and plausibility, we adopt a hybrid strategy for edit-wise explanations, where each edit is explained through three aspects, including error type labels, error severity levels, and free-text error descriptions. 1) Error type labels allow language learners to comprehend and infer syntax and grammar rules in an inductive manner. In particular, we employ a two-tier hierarchical taxonomy including 5 major types and 16 minor types shown in Table 1, inspired by existing studies Liping (2014); Peng et al. (2021); Zhang et al. (2022). The detailed descriptions of various error types are included in Appendix A.1 and A.2. If an edit covers multiple error types, we select the one with the highest granule. 2) Error severity levels, ranging from 1 to 5 points, indicate the significance of a specific grammatical error. 3) Error descriptions are the most crucial and flexible element. These provide keywords, pertinent linguistic knowledge, causes of errors, and revision guidance in a free-text format. We stipulate well-defined error descriptions should meet three principles: fluency, reasonability (making sense to humans), and faithfulness (targeted to a specific edit). To ensure reasonability and faithfulness, the error description must mostly conform to the syllogism form of deductive reasoning: [major premise: semantic rules and related knowledge], [minor premise: the reason for the error in the text], and [explain how to correct it]. Further, any evidence from the source $X$ must be enclosed within special markers 【 】. Similarly, correction content that occurs in the target sentence $Y$ must be enclosed within { }, as indicated in Figure 1.
Major Type Minor Type
Punctuation-level Error 标点冗余 (Punctuation Redundancy)
标点丢失 (Punctuation Missing)
标点误用 (Punctuation Misuse)
Spelling-level Error 字音混淆错误
Phonetic Confusion Error
字形混淆错误
Glyph Confusion Error
词内部字符异位错误
Internal Character Misplacement Error
命名实体拼写错误
Named Entity Misspelling
Word-level Error 词语冗余 (Word Redundancy)
词语丢失 (Word Missing)
词语误用 (Word Misuse)
Sentence-level Error 词序不当 (Improper Word Order)
逻辑不通 (Illogicality)
句式杂糅 (Run-on Sentence)
Other Special Error 照应错误 (Inconsistency Error)
歧义错误 (Ambiguity Error)
语气不协调 (Inconsistent Tone)
Other
Table 1: Hierarchical taxonomy of grammatical error types defined in our benchmark.
4.2 Explanation Synthesizing
Annotating high-quality explanations on a large scale poses a huge challenge to our benchmark construction. Hence, we leverage GPT-4 to synthesize edit-wise explanations efficiently. To achieve this, we first select 10,000 parallel samples across 6 existing benchmarks or datasets of Chinese GEC, with the details listed in Table 2. We pick out only the samples with changed target sentences and select the single target sentence with the most edits if a sample is annotated with multiple target sentences. Then, we prompt GPT-4 to generate edit-wise explanations following in-context learning. To ensure the faithfulness of the synthesized explanation, we first extract edits using the toolkit CLEME Ye et al. (2023c). Inspired by Li et al. (2022a), we then employ the Rationalization Prompting (RP) strategy, where we concatenate task definition, demonstrations, and a parallel sample $(X,Y)$ with extracted edits $E=\{e_{1},e_{2},\cdots,e_{n}\}$ as the prompt. For each error type, we provide the definition, a suggested template of error description, and a demonstration. The prompt is listed in Appendix A.3.
Dataset Sentences Edits/Sent. Chars/Sent.
FCGEC 41,340 1.0 53.1
YACLC-minimal-dev 1,839 2.9 25.9
MuCGEC-dev 1,137 3.2 38.5
NaCGEC-dev 500 1.1 56.2
NLPCC-test 2,000 2.0 29.7
HSK 156,870 1.4 27.2
EXCGEC (FCGEC) 2,308 1.1 55.1
EXCGEC (YACLC) 1,235 3.5 24.3
EXCGEC (MuCGEC-dev) 789 3.3 40.4
EXCGEC (NaCGEC-dev) 449 1.1 56.1
EXCGEC (NLPCC-test) 1,611 1.7 28.9
EXCGEC (HSK) 1,824 2.1 32.0
EXCGEC-train 5,966 2.0 38.7
EXCGEC-dev 750 2.0 38.9
EXCGEC-test 1,500 2.0 39.2
[…]
Table 2: Dataset statistics of the EXCGEC benchmark.
4.3 Explanation Refinement and Analysis
Benefiting from the extensive knowledge acquired during the large-scale pre-training process, GPT-4 is able to generate fluent, reasonable, and plausible explanations in most cases, meeting the requirements with specified instructions. However, GPT-4 is not guaranteed to produce all high-quality explanations due to hallucination, and the patterns of those invalid explanations are referred to as failure modes. Therefore, we hire 12 native speakers, all of whom are graduated students, to screen out invalid explanations. We finally obtain 8,216 clean samples out of 10,000 samples. We further investigate the failure modes of invalid explanations generated by GPT-4, which is provided in Appendix A.4.
4.4 Automatic Metrics
Recent studies leverage human evaluation for the evaluation of GEE due to the lack of enough annotated samples, posing a challenge for the efficient development of EXGEC systems. In this paper, we introduce a comprehensive set of automatic metrics for both correction and explanation parts.
Correction.
We employ CLEME and ChERRANT to evaluate the correction performance. Both are edit-based metrics that output P/R/F0.5 scores, and they have been proven reliable metrics for GEC on CoNLL-2014 Ye et al. (2023c).
Explanation.
Since an edit-wise explanation consists of three critical elements, we define respectively automatic metrics for them. 1) Accuracy and Macro-F1 scores are computed for error type clarification, following the conventional evaluation protocol of text clarification Li et al. (2020). 2) We report the mean absolute error (MAE) to show the deviation of hypothesis error severity levels towards ground truth ones. 3) We employ various metrics for evaluating the free-text explanation description, including BLEU, METEOR, ROUGE-1, ROUGE-2, and ROUGE-L. We leave the analysis on the effectiveness of these metrics to Section 7.2.
5 Method
5.1 Training
To streamline the training process covering all the tasks mentioned in Section 3, we treat all of them as a unified Seq2Seq task Yang et al. (2022b, a). To achieve this, we linearize the data in the format of json Gao et al. (2023). This structured approach simplifies the process of output parsing involving three types elements of edit-wise explanations, and provides a consistent and controllable view to distinguish tasks, enabling the model to understand essential task elements and their relations. Therefore, we train all models using the same smooth cross-entropy loss, regardless of the specific task.
5.2 Inference
Algorithm 1 COTE Decoding Algorithm
1:Source text $X$, a post-explaining model $\mathcal{M}$, and the edit extraction function $f$.
2:Target text $\hat{Y}$, and explanations $\hat{E}^{\prime}$.
3:$\hat{Y}\leftarrow\operatorname{BeamSearch}(\mathcal{M}(\operatorname{Json}(X)))$
4:$\hat{E}^{\prime}\leftarrow\emptyset$
5:if $\hat{Y}=X$ then
6:     return $\hat{Y},\hat{E}^{\prime}$
7:end if
8:$E\leftarrow f(X,\hat{Y})$
9:$\hat{E}^{\prime}\leftarrow\operatorname{Top-P}(\mathcal{M}(\operatorname{Json}(X,Y,E)))$
10:return $\hat{Y},\hat{E}^{\prime}$
For post-explaining EXGEC models, we design a specific Correct-Then-Explain decoding algorithm called COTE, which is presented in Algorithm 1. First, we employ the greedy beam search decoding strategy for the correction part, which is beneficial to relieve the over-correction problem that is common on LLMs Cao et al. (2023); Loem et al. (2023); Li et al. (2023a). Then, we apply CLEME to extract edits. Notably, we merge adjacent edits with distance less than 2 characters to avoid fragmented edits. Finally, we leverage the Top-p decoding strategy for generating explanations, encouraging diversified natural language explanations. It is worth noting that COTE is not accessible to pre-explaining models since the edit extraction tool necessitates both a source text and a target text.
6 Experiments
[…]
Qwen1.5-7B-base 26.00 / 26.54 / 26.10 33.87 / 20.16 / 29.81 67.29 56.81 60.99 29.82 0.80 15.22 39.05 49.74 / 23.28 / 34.32
Qwen1.5-7B-chat 28.31 / 21.21 / 26.54 36.74 / 17.26 / 29.98 68.94 64.83 61.98 29.62 0.75 15.49 38.88 50.32 / 24.25 / 35.24
[…]
Llama3-8B-instruct 21.33 / 26.05 / 22.14 29.00 / 19.40 / 26.39 61.40 55.71 59.16 25.63 0.88 14.70 36.89 49.41 / 23.54 / 34.87
[…]
DeepSeek-7B-chat 25.46 / 18.51 / 23.68 34.02 / 15.75 / 27.62 67.52 66.64 58.11 24.45 0.84 13.94 36.97 48.66 / 22.70 / 34.23
Qwen1.5-7B-chat 13.76 / 13.42 / 13.69 19.27 / 9.93 / 16.22 29.49 80.24 23.35 8.22 1.17 7.75 27.67 40.47 / 15.00 / 28.20
Llama3-8B-instruct 7.12 / 11.17 / 7.68 10.86 / 8.57 / 10.31 23.88 73.06 24.31 8.78 1.21 5.78 23.07 37.57 / 13.47 / 27.19
DeepSeek-7B-chat 9.93 / 8.26 / 9.55 14.28 / 7.07 / 11.86 24.72 78.67 19.12 5.84 1.29 5.91 23.95 37.59 / 13.11 / 26.78
Table 3: Main results of multi-task learning models. Results of post-explaining models are listed in the top block, while those of pre-explaining models are in the bottom block.
[…]
Qwen1.5-7B-base —- / —- / —- —- / —- / —- 98.42 6.14 85.00 43.32 0.73 19.70 43.18 53.48 / 27.79 / 38.12
Qwen1.5-7B-chat 62.59 / 87.35 / 66.35 67.58 / 69.53 / 67.96 99.93 0.43 81.53 39.56 0.73 17.88 41.40 51.73 / 28.81 / 36.51
[…]
Llama3-8B-instruct 69.10 / 90.90 / 72.58 73.75 / 74.37 / 73.87 99.63 1.67 85.99 41.84 0.78 20.73 42.98 54.60 / 29.64 / 40.04
[…]
DeepSeek-7B-chat 41.12 / 79.02 / 45.48 48.35 / 53.20 / 49.25 99.93 0.40 81.17 35.93 0.74 19.57 42.32 53.12 / 28.03 / 38.59
Table 4: Ground truth results of multi-task learning models. We report the explanation performance (right block) of post-explaining models conditioned on source texts and ground truth target texts. Contrarily, we report the correction performance (left block) of pre-explaining models conditioned on source sentences and ground truth explanations.
6.1 Experimental Settings
Backbones.
We benchmark three series of LLMs, including Llama-3 Touvron et al. (2023), Qwen-1.5 Bai et al. (2023), and DeepSeek Bi et al. (2024). For each series of LLMs, we experiment with their base and chat (or instruct) versions to investigate whether further alignment training benefits the task. All results are based on EXCGEC-test. Training details are reported in Appendix B.1.
Evaluation.
We report experiment results using the metrics introduced in Section 4.4, calculated using open-source toolkits including NLTK Bird and Loper (2004), rouge Lin (2004), and scikit-learn Pedregosa et al. (2011). Particularly, we observe many hypothesis edits are not covered in references, making it impossible to evaluate the subsequent explanations for these edits. To address this, we introduce two extra indicators, namely Hit and Miss rates. A hypothesis edit overlapping with a reference edit is designated as a hit edit, while a reference edit without any match with hypothesis edits is deemed a miss edit. The hit rate is defined as the ratio of hit edits to all hypothesis edits, and the miss rate as the ratio of miss edits to all reference edits. Only hit edits are used to determine the evaluation outcomes for explanations.
[…]
Post-explaining 28.31 / 21.21 / 26.54 36.74 / 17.26 / 29.98 68.94 64.83 61.98 29.62 0.75 15.49 38.88 50.32 / 24.25 / 35.24
Pre-explaining 13.76 / 13.42 / 13.69 19.27 / 9.93 / 16.22 29.49 80.24 23.35 8.22 1.17 7.75 27.67 40.47 / 15.00 / 28.20
Pipeline 32.45 / 23.93 / 30.29 40.50 / 19.58 / 33.37 72.00 63.10 65.76 32.77 0.70 16.41 40.04 51.07 / 24.92 / 35.89
Table 5: Comparison of the multi-task solutions and the GEC-GEE pipeline solution based on Qwen1.5-7B-chat.
6.2 Results of Multi-task Models
The preliminary results from both the post- and pre-explaining models are presented in Table 3, from which we can make some conclusions.
Post-explaining models consistently outperform pre-explaining models.
In relation to the correction aspect, all post-explaining models obtain higher F0.5 scores than pre-explaining models, regardless of the applied backbones. A similar pattern is observed in the explanation part, where all the pre-explaining models invariably underperform their post-explaining counterparts. This suggests that a complexity for LLMs in initially explaining grammatical errors. And once pre-explaining models generate flawed explanations, the ensuing distraction impedes their ability to accurately correct the source text.
Chat models outperform base models.
For post-explaining models, we observe all chat or instruct models gain slightly higher F0.5 correction scores, and they also marginally outperform their base version counterparts in the explanation task. It indicates that additional alignment training Wang et al. (2023) can benefit the EXGEC task.
[…]
Beam search 28.31 / 21.21 / 26.54 36.74 / 17.26 / 29.98 99.22 19.05 83.93 44.48 0.71 22.71 44.28 55.55 / 32.26 / 42.34
Top-p 19.45 / 27.05 / 20.61 24.83 / 19.14 / 23.44 99.93 0.40 81.53 39.56 0.74 17.88 41.40 51.73 / 25.81 / 36.51
Table 6: Comparison of the post-explaining model with different token-wise decoding strategies. Note that the explanation performance is conditioned on ground truth target texts in order to exclude unrelated interference.
6.3 Ground Truth Results
In order to study the isolated performance of multi-task models, we provide part ground truth information in advance during the inference stage. Specifically, we provide ground truth target texts for post-explaining and report their performance of explanation. Conversely, we offer ground truth explanations for pre-explaining and report their performance of correction. This experimental setting allows for revealing the specialized performance, eliminating the distraction of previously generated contents. The results are presented in Table 4.
For the task of explanation, two base models slightly outperform chat models.
Specially, the base version models of Qwen and DeepSeek exhibit a minor increase in performance over their chat/instruct counterparts on classifying error types and providing error descriptions. However, this is not true for Llama3, where the Llama3-instruct model obtain the highest Acc, METEOR and ROUGE scores. Also noteworthy is the significantly lower miss rates of chat/instruct models compared to base models, indicating a tendency for the latter to overlook explanations, even when ground truth target texts are available. These findings contradict the joint results in Table 3. We speculate the reason is base models may be more susceptible to low-quality self-generated corrections.
Ground truth explanations tremendously improve correction performance.
Since the explanations include explicit clues for corrections such as evidence words and revision advice, it is effortless for pre-explaining models to correct the source.
6.4 Comparison with Pipeline
We compare the results of multi-task models and GEC-GEE pipeline with COTE in Table 5. It indicates that the pipeline can improve both correction and explanation performance compared to multi-task models, highlighting the challenges of learning a multi-task model for EXGEC. A significant advantage of the pipeline solution is its superior performance of the correction task, which can subsequently enhance the explanation performance. However, adopting the pipeline solution requires heavy deployment and extra training costs.
7 Analysis
7.1 Ablation Results
We conduct ablation studies on Qwen1.5-7B-chat to provide in-depth insights into post-explaining models. We also study the effect of model sizes in Appendix B.2 and provide a case study for different LLMs in Appendix B.3.
[…]
w COTE 99.93 0.43 81.53 39.56 0.74 17.88 41.40 51.73 / 25.81 / 36.51
w/o COTE 49.64 54.01 42.51 17.77 0.93 11.53 33.81 46.35 / 19.34 / 31.28
Table 7: Ablation results of COTE from the same Qwen1.5-7B-chat post-explaining model.
Pearson Spearson
Human v.s. BLEU 0.9222 0.6571
Human v.s. METEOR 0.9280 0.7714
Human v.s. ROUGE-1 0.9464 0.8286
Human v.s. ROUGE-2 0.9175 0.4857
Human v.s. ROUGE-L 0.9352 0.6571
A1 v.s. A2 0.9874 0.9429
Table 8: Correlations between human judgements and metrics for error descriptions.
Effect of COTE.
The impact of COTE introduced in Section 5.2 is examined in this section. We provide the post-explaining model with ground truth target texts, which allows us to focus on the explanation performance. The results presented in Table 7 reveal a huge performance drop if we do not leverage COTE, especially the hit and miss rates. This demonstrates the effectiveness of COTE.
Effect of token-wise decoding strategies.
By default, we employ beam search decoding for corrections and top-p decoding for explanations. In this section, we explore the reverse setting, and the results are reported in Table 6. When switching from beam search to top-p for correction, we observe a huge performance drop in precision and F0.5 and increase in recall, which means top-p encourages LLMs to over-correct Cao et al. (2023). On the other hand, leveraging beam search improves explanation performance, suggesting the potential benefits of a greedy decoding algorithm for the task. However, we notice that beam search also increases the miss rate. We speculate that beam search may discard some low-likelihood explanations.
7.2 Human Evaluation for Error Descriptions
Despite the efficiency of automatic metrics in evaluating error descriptions, their accuracy remains to be confirmed. Therefore, this section attempts to demonstrate the suitability of different metrics by comparing their corrections with human judgements. We report the correlations between two human annotators and the ones between average human ratings and metric scores in Table 8. We observe the inter-annotator correlations are close to 1, meaning it is relatively easy to determine the quality of error descriptions for human. Among various metrics, ROUGE-1 achieve the highest correlations, followed by METEOR. All the introduced metrics show moderate or high correlations, indicating that it is advisable to employ them as proxies for human evaluation. We provide the detailed annotation guidance and rating rules in Appendix B.4.
8 Conclusion
We propose and formulate the task of EXGEC, overcoming the limitation of previous studies that fail to establish the interaction of both correction and explanation tasks. To develop the task, we propose the EXCGEC benchmark, based on which we develop baseline models in multiple settings. Extensive experiments and analyses reveal several challenges of the task, and we hope this paper can serve as a starting point for future exploration.
Limitations
Inferior performance of multi-task models.
In our experiments, we observe the pipeline solution outperform the multi-task solutions, regardless of correction or explanation tasks. This suggests that the multi-task models struggle to reap positive benefits from the interaction of both tasks. We leave the exploration of effective multi-task learning EXGEC models to the future work.
Limitations of synthesizing datasets.
LLM-augmented datasets may include some unintended biases towards or inaccuracies, resulting in skewed or unfair outcomes in applications. Second, it is necessary to manually filter out invalid data in order to ensure the quality of datasets. But it is indeed a advisable method to construct datasets using LLMs, considering its efficiency.
Adaptation to other languages.
The general design of our proposed edit-wise explanations can be easily adapted to other languages. However, the detailed design may not be suitable to other languages. For example, the two-tier hierarchical taxonomy of error types is tailored for Chinese.
Ethics Statement
Our proposed benchmark is built upon existing datasets, backbones and metrics, all of which are publicly available. We have cited the corresponding authors or projects of them, and confirm that they are consistent with their intended use.
Additionally, we conduct human evaluation experiments to ensure the quality of the dataset and find out the correlations between metrics and human judgements. To achieve this, we hire 12 native speakers, all of whom are graduated students. Each annotator could complete the entire annotation process within approximately 6$\sim$8 working hours. All annotators were paid for their work, with an average salary of approximately \$5 per hour.
Appendix A Benchmark Details
A.1 Description of Grammatical Error Types
In the taxonomy of Chinese grammatical errors, we first divide grammatical errors by their effecting granular into 5 major types, including punctuation-level, spelling-level, word-level, sentence-level, and other special errors. This section will clarify all involved grammatical error types by providing detailed descriptions and corresponding examples listed in Figure 456.
Punctuation-level Error.
This type of grammatical errors primarily involves redundancy, missing, and misuse of punctuation.
  • 标点冗余 (Punctuation Redundancy). The type of punctuation redundancy refers to the insertion of punctuation in unnecessary places. For punctuation redundancy errors, first explain the role of the punctuation symbols involved, and then explain the reasons for punctuation redundancy in the current case.
  • 标点丢失 (Punctuation Missing). Punctuation missing mainly refers to the omission of punctuation that should have existed in the middle and end of a sentence. For the explanation of punctuation missing errors, first point out the evidence words and missing punctuation symbols, and then explain the role of the added punctuation in this context.
  • 标点误用 (Punctuation Misuse). Misuse of punctuation is very common in daily Chinese writing. For cases of punctuation misuse, first briefly explain the roles of misused punctuation and correct punctuation, and then explain sufficient reasons for correction.
Spelling-level Error.
Spelling-level errors refer to people who, due to carelessness or lack of knowledge, write incorrect characters or words during the writing process. The type is so common that Chinese Spelling Check (CSC), as a standard NLP task specialized in spelling-level errors, attracts the attention of many researchers. Inspired by these studies, we categorize spelling-level errors further into 4 sub-classes.
  • 字音混淆错误 (Phonetic Confusion Error). Phonetic confusion errors are caused by misusing Chinese characters with the same or similar pinyin. The vast majority of Chinese Internet users use the pinyin input method, so many Chinese spelling-level errors on the Internet fall into this type.
  • 字形混淆错误 (Glyph Confusion Error). In addition to the pinyin input method, some users apply the Wubi input method or other glyph-based input methods. In this case, they are prone to spelling errors due to confusion of fonts or strokes.
  • 词内部字符异位错误 (Internal Character Misplacement Error). Internal character misplacement error refers to expressing a multi-character word in the disorder of characters. The type seldom happens for native speakers, but sometimes in texts written by L2 speakers. For example, the spelling-level error “共公” falls in this type and should be corrected to “公共”.
  • 命名实体拼写错误 (Named Entity Misspelling). There are numerous named entity words in Chinese, such as person names, organization names, place names, and all other entities identified by terminologies. These words are also very prone to spelling errors.
Word-level Error.
Word-level errors often refer to misuse of individual words or idioms in a sentence, but the syntactic structure of the sentence is correct. This type of error belongs to the most common category in Chinese text errors and can usually be subdivided into the following three types:
  • 词语冗余 (Word Redundancy). The simultaneous appearance of words with the same or similar meanings in a sentence can cause semantic repetition and sentence redundancy, which is known as word redundancy. Repeated words often appear adjacent to each other, so it is important to pay attention to whether the meanings of adjacent words are exactly the same. If they are the same, it may lead to the problem of word redundancy.
  • 词语丢失 (Word Missing). In modern Chinese, sentences generally have six major components, namely subject, predicate, object, attributive, adverbial, complement, etc. A sentence must express a complete meaning, and its structure must also be complete. The so-called complete structure does not mean that a sentence must have the usual six components, but rather that the sentence should be composed of the necessary components to express the complete meaning. If the necessary sentence components are missing, it will cause the phenomenon of word missing.
  • 词语误用 (Word Misuse). Word Misuse indicates improper use of words in the text. The main cause of this error is the author’s insufficient understanding of the meaning and part of speech of a certain word.
Sentence-level Error.
This type mainly involves sentence-level issues, not just individual words or characters. Sentence-level errors are often caused by violating common syntactic structures, or not following objective reasoning.
  • 词序不当 (Improper Word Order). Proper word order is essential to express exact meaning in Chinese. Writing texts without accurate word order results in the type of improper word order. If a sentence is not combined according to the intended meaning, it may lead to confusion in the sentence structure, resulting in an imbalance in the relationship between sentence components and affecting the expression of sentence meaning.
  • 逻辑不通 (Illogicality). Illogicality refers to a sentence that conforms to grammatical norms but does not conform to logical reasoning. Illogicality can be caused by many reasons such as improper logical order, causal confusion, and reversal of subject and object.
  • 句式杂糅 (Run-on Sentence). Run-on Sentence in Chinese usually refers to the use of two formats or sentences with similar or identical meanings in one sentence. People originally used one format when writing sentences, but due to interference from other factors such as sentence content, they may unconsciously switch to another format, resulting in a mixture of the two formats.
Other Special Error.
Besides the above grammatical error types, other several types can not easily fit in the mentioned major types. So we classify them to other special errors.
  • 照应错误 (Inconsistency Error). Inconsistency errors are ones involved in the mistaken referential relationship between two words, and explaining this grammatical error requires knowledge of the referential relationship between each word.
  • 歧义错误 (Ambiguity Error). Ambiguity errors happen when a word or a sentence can be understood as having multiple meanings.
  • 语气不协调 (Inconsistent Tone). Inconsistent tone refers to the inconsistency of tone between the preceding and following sentences.
Additionally, we define the grammatical error type Other as ones that do not fit in any of the above error types. These errors are usually involved in rather significant modification and sometimes change the original semantics.
Figure 3: Distribution of 7 kinds of LLM errors.
Configuration Value
Fine-tuning
Devices 2 Tesla A100 GPU (80GB)
Epochs 5
Finetuning type Lora
Train batch size per GPU 2
Eval batch size per GPU 1
Gradient accumulation steps 16
Optimizer
AdamW
($\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1\times 10^{-6}$)
Learning rate $5\times 10^{-5}$
Learning rate schedule cosine decay
Warmup steps 20
Eval steps 200
Cutoff length 1024
Preprocessing workers number 16
Numerical precision fp16
Weight decay 0.05
Inference
Beam size 5
Top-p 0.8
Max new tokens 2048
Temperature 0.7
Table 9: Hyper-parameters used in our experiments.
[…]
Qwen1.5-1.8B-chat 21.11 / 19.28 / 20.72 28.91 / 15.70 / 24.74 59.94 65.14 55.80 23.27 0.89 10.19 34.35 48.66 / 22.70 / 34.23
Qwen1.5-4B-chat 22.49 / 20.84 / 22.14 30.57 / 16.85 / 26.29 62.91 62.70 57.16 25.31 0.85 11.61 35.91 46.83 / 19.59 / 30.86
Qwen1.5-7B-chat 28.31 / 21.21 / 26.54 36.74 / 17.26 / 29.98 68.94 64.83 61.98 29.62 0.75 15.49 38.88 50.32 / 24.25 / 35.24
Table 10: Comparison of post-explaining models with various model sizes.
Figure 4: Examples of error types.
Figure 5: Examples of error types.
Figure 6: Examples of error types.
A.2 Examples of Error Types
We list the examples of error types in Figure 456.
A.3 Prompt of Generating Explanations
The prompt we use to generate explanations is shown in Figure 8. We also provide an English version in Figure 9.
A.4 Detailed Description of LLM Failure Modes
We categorize the failure modes in our case into seven major reasons: incorrect type, incorrect severity, incorrect format, incorrect template, non-fluency, unreasonability, and unfaithfulness. One expert annotator is asked to classified the sampled 100 invalid explanations, where an explanation may be categorized into multiple failure modes. The annotation results, illustrated in Figure 3, reveal that GPT-4 tend to mis-classify grammatical errors and providing unfailthful error descriptions. On the other hand, GPT-4 is capable to a large extend to offer well formed, fluent, and reasonable explanations, demonstrating the effectiveness of LLM annotation on this task.
The definitions of seven failure modes of explanations are as follows:
  • Incorrect type: the error type is incorrect.
  • Incorrect format: the evidence content and the correction content are not highlighted by special markers 【 】 or { }.
  • Incorrect template: the error description does not follow the syllogism form of deductive reasoning.
  • Non-fluency: the error description is non-fluent or unreadable.
  • Unreasonability: the error description contains obvious mistakes about linguistics, thus making it unacceptable for human.
  • Unfaithfulness: the error description is not targeted to the given edit.
Appendix B Experimental Details and Extra Results
B.1 Implementation Details.
We train all models for 5 epochs and select the best model validated on EXCGEC-dev and report its performance on EXCGEC-test. The detailed training hyperparameter values of the all models in our experiments are shown in Table 9.
B.2 Effect of Model Sizes
Table 10 indicates the varying performance across model sizes ranging from 1.8B to 7B. We observe consistent performance enhancement with increasing model sizes.
B.3 Case Study
We provide a case study in Table 7.
B.4 Details of Human Rating
Specifically, we hire 2 native Chinese speakers to rate the explanations generated by 6 post-explaining models in Table 3 conditioned on ground truth target texts. The rating scores range from 0 to 100, and each annotator concurrently rate 6 explanations for each sample. We randomly select 100 samples for annotation. We provide annotators with general scoring suggestions:
  • 100 points: Explain and describe fluently (fluency), introduce relevant semantic knowledge to enhance persuasiveness (rationality), and explain that it is aimed at the current editor (loyalty). All aspects are impeccable, and there is almost no better explanation or description than this.
  • 80$\sim$100 points: The explanation and description are expressed fluently, satisfy fidelity, and have a certain degree of rationality, but there are certain degrees of flaws.
  • 60$\sim$80 points: The explanation and description are expressed fluently, but the fidelity or rationality is not good enough, but it is somewhat helpful for correcting the grammar error in understanding.
  • 30$\sim$60 points: The explanation and description are expressed fluently, but the rationality is poor, and it is not very helpful for correcting the grammar error in understanding.
  • 0$\sim$30 points: The explanation and description are expressed fluently, but the fidelity is poor, and the object of explanation is not the current editor. There is no help in correcting the grammar error for understanding.
  • 0$\sim$30 points: The explanation and description are vague and cannot be understood. There is no help in correcting the grammar error for understanding.
Figure 7: A case study of all the LLMs involved in our experiments.
Figure 8: The prompt used for explanation generation. For each error type, We provide the definition, a suggested template of error description, and a demonstration for GPT-4.
Figure 9: The English prompt used for explanation generation.